# MCGAN: Enhancing GAN Training with Regression-Based Generator Loss

### v1

@@ 3.4 Discriminability and optimality of MCGAN
Assume Assumption 3.1 holds, and let $\phi^{\prime}_{\cdot,\cdot}:\mathcal{P(X)}\times\mathcal{P(X)}\rightarrow\Phi$ be a parameterization map such that $D^{\phi^{\prime}_{\cdot,\cdot}}:\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})\times\mathcal{X}\rightarrow\mathbb{R}$ has discriminability, i.e. $D^{\phi^{\prime}_{\cdot,\cdot}}\in\mathcal{D}_{\text{Dis}}$. If $\theta^{*}$ is a local minimizer of $\mathcal{L}_{G}(\theta;\phi^{\prime}_{\mu,\nu_{\theta}},\mu)$ defined in (5), then $\nu_{\theta^{*}}=\mu$.
@@
Theorem 3.1 demonstrates that to learn the data distribution $\mu$, MCGAN do not restrict the discriminator to reach its optimum; the discriminability is sufficient, which is again attributed to the strong supervision provided by regression loss $\mathcal{L}_{R}$.
4 Properties of MCGAN
4.1 Relation to $f$-divergence
@@
Given the optimal discriminator (2), the vanilla GAN objective (1) can be interpreted as minimizing Jensen-Shannon divergence between $\mu$ and $\nu_{\theta}$, subtracting a constant term $\log(4)$. The generator is therefore trained to minimize the Jensen-Shannon divergence. Similarly, [8] also showed that optimizing LSGANs yields minimizing the Pearson $\chi^{2}$ divergence between real and fake measures.
@@
Given (13), we are also able to explore the connection between MCGAN and $f$-divergence. As proven in Lemma 4.1, when considering the optimal discriminator (13), the difference between real and fake expected discriminator output in (8) can be interpreted as an $f$-divergence between $\mu$ and $\bar{\nu}$, where $\bar{\nu}:=\frac{(\mu+\nu)}{2}$ represents the averaged measure with density $p_{\bar{\nu}}(x):=\frac{p_{\mu}(x)+p_{\nu_{\theta}}(x)}{2}$.
Lemma 4.1
@@
Lemma 4.1 establishes a connection between MCGAN and $f$-divergence, illustrating the information-theoretic aspects of the MCGAN framework. Unlike the $\text{KL}(\nu|\mu)$ induced in non-saturating loss [6], this $\text{Div}_{f}(\mu|\bar{\nu})$ can avoid mode dropping by assigning moderate cost on the occasions where $p_{\mu}(x)\gg p_{\nu_{\theta}}(x)$.
4.2 Improved stability
Lack of stability is a well-known issue in GAN training, and it arises due to several factors. Arjovsky et al. [6] provides insights into this instability issue of non-saturating loss. The instability is analysed by modelling the inaccurate discriminator as an optimal discriminator perturbed by a centred Gaussian process. Given this noisy version of the optimal discriminator, it can be shown that the gradient of non-saturating loss follows a centred Cauchy distribution with infinite mean and variance, which leads to massive and unpredictable updates of the generator parameter. Hence it can be regarded as the source of training instability.
@@ 4.2 Improved stability
Theorem 4.1
Let $D^{\phi_{\epsilon}}(x)$ be a noisy version of optimal discriminator such that $D^{\phi_{\epsilon}}(x)=D^{\phi^{*}}(x)+\epsilon_{1}(x)$ and $\nabla_{x}D^{\phi_{\epsilon}}(x)=\nabla_{x}D^{\phi^{*}}(x)+\epsilon_{2}(x)$ for $\forall x\in\mathcal{X},$ where $\epsilon_{1}(x)$ and $\epsilon_{2}(x)$ are two uncorrelated and centred random noises that are indexed by $x$ and have finite variance.222 This assumption of centred random noise is made due to the fact that as the approximation gets better, this error looks more and more like centred random noise due to the finite precision [6]. Then for $\mathcal{L}_{G}(\theta;\phi)$ in (5), we have
@@
Theorem 4.1 implies that given an inaccurate discriminator, the expected value of the gradient of $\mathcal{L}_{G}(\theta;\phi)$ in (5) is the accurate gradient given by the optimal discriminator and its variance is finite. More importantly, its variance is determined by the discrepancy between $\mu$ and $\nu_{\theta}$, specifically when the fake measure produces the real measure, the gradient is zero almost surely, indicating improved training stability.
4.3 Relation to feature matching
In order to enhance the generative performance, a feature matching approach is proposed in [25] which adds to the generative loss function an additional cost that matches the statistic of the real and generated samples given by the activation on an intermediate layer of the discriminator. The generator hence is trained to generate fake samples that reflect the statistics (features maps) of real data rather than just maximizing its discriminator outputs.
@@ 4.3 Relation to feature matching
$R_{\text{fm}}(\theta;\phi)=\|\mathbb{E}_{\mu}[\psi(x)]-\mathbb{E}_{\nu_{\theta}}[\psi(x)]\|^{2}_{2}.$ (17)

### v3

@@ Introduction
Assume Assumption 1 holds, and let $\phi^{\prime}_{\cdot,\cdot}:\mathcal{P(X)}\times\mathcal{P(X)}\rightarrow\Phi$ be a parameterization map such that $D^{\phi^{\prime}_{\cdot,\cdot}}:\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})\times\mathcal{X}\rightarrow\mathbb{R}$ has discriminability, i.e. $D^{\phi^{\prime}_{\cdot,\cdot}}\in\mathcal{D}_{\text{Dis}}$. If $\theta^{*}$ is a local minimizer of $\mathcal{L}_{G}(\theta;\phi^{\prime}_{\mu,\nu_{\theta}},\mu)$ defined in (3), then $\nu_{\theta^{*}}=\mu$.
@@
Theorem 1 implies that MCGAN can effectively learn the data distribution $\mu$ without requiring the discriminator to reach its optimum; the discriminability is sufficient, which is again attributed to the strong supervision provided by regression loss $\mathcal{L}_{R}$. We defer the proof of Theorem 1 and other theoretical properties of MCGAN, e.g., improved training stability and relation to $f$-divergence to the Appendix.
Numerical experiments
To validate the efficacy of the proposed MCGAN method, we conduct extensive experiments on a broad range of data, including image, time series, and video data for various generative tasks. For image generation, the conditioning variables are categorical, whereas for time series and video generation tasks, the conditioning variables are continuous. To show the flexibility of MCGAN to enhance different GAN backbones, we choose several state-of-the-art GAN models with different discriminative losses (i.e., BCE and Hinge loss) as baselines. Various test metrics and qualitative analysis are employed to give a comprehensive assessment of the quality of synthetic data generation.
The full implementation details of numerical experiments, including models, test metrics, hyperparameters, optimizer and supplementary numerical results, can be found in Appendix E. Moreover, we will open-source the codes and final checkpoints upon publication for reproducibility.
Unconditional and conditional image generation
Datasets
We conduct conditional image generation tasks using the CIFAR-10 and CIFAR-100 datasets (Alex 2009), which are standard benchmarks with 60K 32x32 RGB images across 10 and 100 classes, respectively. To further validate our MC method on larger and higher-resolution datasets, we include: 1) the unconditional FFHQ256 dataset, which contains 70K 256x256 human face images, 2) the conditional ImageNet64 dataset, which has 1.2 million 64x64 images across 1,000 classes, and 3) the unconditional LSUN bedroom data, which has 3 million 256x256 images.
We validate our method using two different backbones, BigGAN (Brock, Donahue, and Simonyan 2018) and StyleGAN2 (Karras et al. 2020b). The test metrics include Inception Score (IS), Fréchet Inception Distance (FID), and Intra Fréchet Inception Distance (IFID) together with two recognizability metrics Weak Accuracy (WA) and Strong Accuracy (SA). To alleviate the overfitting and improve the generalization, we also increase data efficiency by using the Differentiable Augmentation (DiffAug) (Zhao et al. 2020).
In the following, we mainly focus on the CIFAR-10 dataset for in-depth analysis, with a brief summary of the results on the other datasets.
Faster training convergence
Improved fidelity metrics
As shown in Table 2, our MC method considerably improves all the baselines independently of the choice of discriminative loss ($\mathcal{L}_{D}$). Specifically, when using Hinge loss as $\mathcal{L}_{D}$ along with DiffAug, the MC method improves the FID from 4.43 to 3.61, comparable to the state-of-the-art FID result of (Kang, Shin, and Park 2022). Also, its IS score is significantly increased from 9.61 to 9.96, indicating better diversity of the generated samples.
In addition, applying the MC method to the cStyleGAN2 backbone results in an FID improvement of approximately 0.08. Notably, the combination of Hinge + MC + DiffAug achieves an FID of 2.16, which, to our knowledge, is the best FID achieved using StyleGAN2 as the backbone (Kang et al. 2021; Kang, Shin, and Park 2022; Tseng et al. 2021)
Loss Hinge BCE
BigGAN 9.27 $\pm$ 0.11 5.31 16.20 9.30 $\pm$ 0.14 5.55 16.62
   +DiffAug 9.61 $\pm$ 0.06 4.43 14.60 9.51 $\pm$ 0.11 4.71 14.83
   +MC 9.66 $\pm$ 0.09 4.51 14.71 9.62 $\pm$ 0.09 4.61 14.82
   +MC+DiffAug 9.96 $\pm$ 0.12 3.61 13.60 9.94 $\pm$ 0.10 3.93 13.72
StyleGAN2 - - - 10.17 $\pm$ 0.12 3.7 14.04
   +DiffAug 10.19 $\pm$ 0.11 2.25 11.40 10.03 $\pm$ 0.09 2.44 11.62
   +MC+DiffAug 10.26 $\pm$ 0.08 2.16 11.04 10.10 $\pm$ 0.11 2.36 11.30
Table 2: Quantitative results of image generation on CIFAR-10 using BigGAN/StyleGAN2 w/o and with our MC method and Differentiable Augmentation.
Improved recognizability metrics
We generated 10k (the same setting as the test set) images using the BigGAN backbone. The WA rates are 62.56%, 52.09%, and 54.71% for the real test set, the generated set from Hinge baseline, and the generated set from Hinge + MC, respectively. Our MC method’s images perform closer to the real test set than the baseline’s, showing better distribution matching to the real data in terms of recognizability. The SA rate of our MC method is 83.42% compared to 93.65% of the real test set, showing that we generate fairly recognizable fake images.
Qualitative results
The qualitative results are shown in Figure 3 and Figure 9 in Appendix with only a small amount of images (in red boxes) misclassified by our classifier.
Loss Hinge BCE
BigGAN 10.73 $\pm$0.10 8.31 83.36 10.81 $\pm$0.14 8.37 81.89
   +DiffAug 10.72 $\pm$0.13 7.37 80.00 10.71 $\pm$0.08 7.61 80.48
   +MC 11.39 $\pm$0.10 6.97 80.20 11.59 $\pm$0.12 6.99 80.91
   +MC+DiffAug 11.81 $\pm$ 0.06 5.77 76.26 11.90 $\pm$0.08 5.85 77.33
Latent space analysis
The latent space learned by the generator is expected to be continuous and smooth so that small perturbations on the conditional input can lead to smooth and meaningful modifications on the generated output. To explore the latent space, we interpolate between each pair of randomly generated images by linearly interpolating their conditional inputs. The results are shown in Figure 4. Intermediary images between a pair of images from two different classes are shown in each row with their confidence score distributions below. The labels of the two classes are shown on the left and right sides of each row, respectively. Each distribution of the confidence scores is calculated by the bottleneck representation of the ResNet-50 classifier with a softened softmax function of temperature 5.0 for normalization. The score bars of the left class and the right class are shown in green and magenta, respectively. The red boxes highlight the images being classified as a third class, while the yellow boxes highlight the images having non-monotonic transitions of their confidence scores compared to those of their adjacent images. In other words, images in both red and yellow boxes are undesirable as they imply that the latent space is less continuous and less smooth. By comparing Figure 4a and 4b, we can see that the MC method outperforms in the learned latent space and has most of the decision switch between the two classes occur in the middle range of the interpolation.
Quantitative results on CIFAR-100
For completeness, we show the image generation performance on CIFAR-100 in Table 3. Significant improvements are achieved by using our MC method independently for both baseline discriminative losses, with an average improvement of 1.1 in IS, 1.6 in FID, and 3.7 in IFID. A detailed sensitive analysis w.r.t the Monte Carlo sample size is provided in the appendix.
Large-scale and high-resolution dataset results
For the FFHQ256 (high-resolution), the lmageNet64 (large-scale), and the LSUN bedroom (large-scale high-resolution) dataset, we use the StyleGAN2-ada (Karras et al. 2020a) as backbones. As shown in Table 4222Baseline results differ from StyleGAN2-ADA’s official benchmarks due to hyperparameter adjustments for different GPU setups., MCGAN achieved significant and consistent gains in both FID and IS, as evidenced by 16.4% ($4.51\rightarrow 3.77$), 15.5% ($19.83\rightarrow 16.76$), and 35.7% ($4.34\rightarrow 2.79$) FID improvement, respectively, on FFHQ256, ImageNet64, and LSUN bedroom datasets. These improvements are significant and consistent during training periods and across various datasets, demonstrating faster convergence and better generation ability.
[…]
FFHQ256 original 4.51 $\pm$ 0.03 5.10 $\pm$ 0.07 0.69 0.40
   +MC 3.77 $\pm$ 0.04 5.25 $\pm$ 0.06 0.69 0.45
ImageNet64 original 19.83 $\pm$ 0.02 13.67 $\pm$ 0.17 0.65 0.33
   +MC 16.76 $\pm$ 0.08 13.96 $\pm$ 0.22 0.63 0.43
LSUN bedroom original 4.34 $\pm$ 0.03 2.45 $\pm$ 0.02 0.57 0.22
   +MC 2.79 $\pm$ 0.01 2.45 $\pm$ 0.02 0.61 0.23
Table 4: Quantitative results of image generation on large-scale and high-resolution datasets using StyleGAN2-ada w/o and with our MC method; FID is 10-run average.
Conditional video generation
The conditional video generation task aims to generate the next frame given the past frames of the videos. Here, we used the Moving MNIST data set (Srivastava, Mansimov, and Salakhudinov 2015), which consists of 10,000 20-frame 64x64 videos of moving digits. The whole dataset is divided into the training set (9,000 samples) and the test set (1,000 samples). For the architecture of both the generator and discriminator, we use the convolutional LSTM (ConvLSTM) unit proposed by (Shi et al. 2015) due to its effectiveness in video prediction tasks. In the model training, the generator takes in 5 past frames as the input and generates the corresponding 1-step future frame, then the real past frames and the generated future frames are concatenated along time dimension and put into the discriminator.
For comparison, we used classical GAN as the benchmark. We trained our model for 20,000 epochs with batch size 16. The model performance is evaluated by computing the MSE between the generated frames and the corresponding ground truth on the test set. Numerical results show that our proposed MC method reduces GAN’s MSE from 0.1012 to 0.0840. Compared to the baseline, the predicted frames from our MC method are clearer, more coherent, and visually closer to the ground truth, as shown in Figure 5.
Figure 5: Results of predicting the next frame given the past 5 frames using ConvLSTM w/o and with our MC method.
Conditional time-series generation
Following (Liao et al. 2024), we consider the conditional time-series generation task on two types of datasets (1) $d$-dimensional vector auto-regressive (VAR) data and (2) empirical stock data. The goal is to generate 3-step future paths based on the 3-lagged value of time series. We apply the MCGAN to the RCGAN baseline (Esteban, Hyland, and Rätsch 2017) and bencharmak it with TimeGAN (Yoon, Jarrett, and Van der Schaar 2019), GMMN (Li, Swersky, and Zemel 2015) and SigWGAN (Liao et al. 2024) as the strong SOTA models for conditional time series generation. The model performance is evaluated using metrics in (Liao et al. 2024) including (1) ABS metric, (2) Correlation metric, (3)ACF metric and (4) $R^{2}$ error to assess the fitting of synthetic data in terms of marginal distribution, correlation, autocorrelation and usefulness, respectively.
VAR dataset
To validate the performance of MCGAN for multivariate time series systematically, we apply our method on VAR datasets with various path dimensions $d\in[1,100]$ and different parameter settings. For $d\in\{1,2,3\}$, MCGAN consistently outperforms the RCGAN and TimeGAN (see Table 9 - 11 in the Appendix). Figure 6 shows that the MCGAN and SigCWGAN have a better fitting than other baselines in terms of conditional law as the estimated mean (and standard deviation) is closer to that of the ground truth compared with the other baselines for $d=3$. Note that SigCWGAN suffers the curse of dimensionality resulting from large $d$ and becomes infeasible for $d\geq 50$, whereas MCGAN does not. In fact, as $d$ increases, the performance gains of MCGAN become more pronounced, which is shown in Table 5. For example, with $d=100$, the MC method improves all the metrics by 30%-40%, further highlighting its effectiveness in high-dimensional settings.
Loss Hinge BCE
$d$ Method ABS $\downarrow$ Corr $\downarrow$ ACF $\downarrow$ ABS $\downarrow$ Corr $\downarrow$ ACF $\downarrow$
10 RCGAN 0.01802 0.05678 0.08175 0.01525 0.05074 0.08995
   +MC 0.01550 0.04361 0.06511 0.01393 0.04589 0.07193
50 RCGAN 0.03527 0.07000 0.08844 0.03632 0.07103 0.08765
   +MC 0.02861 0.06161 0.06869 0.02503 0.05995 0.06996
100 RCGAN 0.03319 0.07895 0.11018 0.03788 0.07295 0.10222
   +MC 0.02296 0.04981 0.06687 0.02344 0.05024 0.06143
Table 5: Quantitative results of time-series generation on VAR data with different path dimensions $d$ ranging from $10$ to $100$ using RCGAN w/o and with our MC method.
Figure 6: Comparison of all models’ performance in fitting the conditional distribution of future time series given one past path sample. The real and generated paths are plotted in red and blue, respectively, with the shaded area as the $95\%$ confidence interval. The training dataset is synthesized from VAR(1) model for $d=3$, $\phi=0.8$ and $\sigma=0.8$.
Stock dataset
The stock dataset is a 4-dimensional time series composed of the log return and log volatility data of S&\&&P 500 and DJI spanning from 2005/01/01 to 2020/01/01. To cover the stylized facts of financial time series like leverage effect and volatility clustering, we also evaluate our generated samples using the ACF metric on the absolute return and squared return. Table 6 demonstrates that our MC method consistently improves the generator performance in terms of temporal dependency, cross-correlation and usefulness. Although RCGAN achieved comparable ABS metrics, it failed to capture the cross-correlation and temporal dependence. Specifically, using our proposed MC method, the correlation metric and ACF metric of RCGAN can be improved from 0.25184 to 0.15687 and from 0.03814 to 0.02905. The gap in the $R^{2}$ further showcases that our MC method can enhance the generator to generate high-fidelity samples.
Model ABS $\downarrow$ ACF $\downarrow$ ACF($|x|$) $\downarrow$ ACF($x^{2}$) $\downarrow$ Corr $\downarrow$ $R^{2}$ (%) $\downarrow$
RCGAN 0.00868 0.03814 0.07874 0.13933 0.25184 4.49683
MCGAN (ours) 0.00996 0.02905 0.05437 0.09933 0.15687 2.84285
SigCWGAN 0.00960 0.02982 0.13385 0.08456 0.11721 3.81981
GMMN 0.01389 0.05989 0.25295 0.26960 0.31838 11.87578
TimeGAN 0.01100 0.05716 0.06899 0.12584 0.47344 4.53960
Table 6: Quantitative results of time-series generation on SPX/DJI data using RNN w/o and with our MC method.
Conclusion
This paper presents a general MCGAN method to tackle the training instability, a key bottleneck of GANs. Our method enhances generator training by introducing a novel regression loss for (conditional) GANs. We establish the optimality and discriminability of MCGAN, and prove that the convergence of optimal generator can be achieved under a weaker condition of the discriminator due to the strong supervision of the regression loss. Moreover, extensive numerical results on various datasets, including image, time series data, and video data, are provided to validate the effectiveness and flexibility of our proposed MCGAN and consistent improvements over the benchmarking GAN models.
For future work, it is worthwhile to explore the application of MCGAN to enhance state-of-the-art GAN models for more challenging and complex tasks, such as text-to-image generation. Moreover, given the flexibility and promising results of the MCGAN on different types of data, it can be effectively applied to generate multi-modality datasets simultaneously.
Acknowledgements
HN and WY are supported by the EPSRC under the program grant EP/S026347/1 and the Alan Turing Institute under the EPSRC grant EP/N510129/1. WY is also supported by the Data Centric Engineering Programme (under the Lloyd’s Register Foundation, UK grant G0095). The authors are grateful to Richard Hoyle from UCL, Terry Lyons from Oxford, and the Oxford Mathematical Institute IT staff for their support with computing resources. We also thank Lei Jiang for his help with some of the numerical experiments and the anonymous referees for their constructive suggestions, which significantly improved the paper.
Appendix A Algorithm
The pseudocode of the MCGAN is provided in Algorithm 1.
Algorithm 1 Algorithm of MCGAN
1:Input:
2:      $N$: number of epochs;
3:      $N_{D}$: number of discriminator iterations per generator iteration;
4:      $B\in\mathbb{N}$: batch size;
5:      $N_{MC}$: number of Monte Carlo samples;
6:      $f_{1},f_{2}$: specified functions used to compute discriminative loss;
7:Output:
8:      $(\theta^{*},\psi^{*})$: approximation of the optimal parameters of the generator and discriminator.
9:Initialize model parameters $(\theta,\psi)$ for generator $G$ and discriminator $D$.
10:for $n=1$ to $N$ do
11:     for $n_{d}=1$ to $N_{D}$ do
12:         Sample batch $\{(x_{i},y_{i})\}_{i=1}^{B}\sim p_{d}(X,Y)$
13:         Generate samples $\{(\hat{x}_{i},y_{i})\}_{i=1}^{B}\sim p_{\theta}(X,Y)$
14:         Compute discriminative loss:
$\mathcal{L}_{D}(\psi;\mu,\nu_{\theta})\leftarrow$ $\frac{1}{B}\sum_{i=1}^{B}f_{1}(D^{\psi}(x_{i},y_{i}))$
$+\frac{1}{B}\sum_{i=1}^{B}f_{2}(D^{\psi}(\hat{x}_{i},y_{i}))$
15:         Update discriminator parameters:
$\psi\leftarrow\text{Adam}(\mathcal{L}_{D})$
16:     end for
17:     Sample batch $\{(x_{i},y_{i})\}_{i=1}^{B}\sim p_{d}(X,Y)$;
18:     For each label $y_{i}$, estimate the conditional expectation:
$\hat{\mathbb{E}}_{p_{\theta}}[D^{\psi}(X,y_{i})\mid y_{i}]\leftarrow\frac{1}{N_{MC}}\sum_{j=1}^{N_{MC}}D^{\psi}(G^{\theta}(y_{i},z^{(j)}),y_{i});$
19:     Compute generative loss:
$\mathcal{L}_{G}\leftarrow$ $\frac{1}{B}\sum_{i=1}^{B}\|D^{\psi}(x_{i},y_{i})-\hat{\mathbb{E}}_{p_{\theta}}[D^{\psi}(X,y_{i})\mid y_{i}]\|^{2};$
20:     Update generator parameters:
$\theta\leftarrow\text{Adam}(\mathcal{L}_{G});$
21:end for
Appendix B Properties of MCGAN
In this section, we explore some favorable properties of MCGAN to further support why it can achieve such advantageous numerical results.
Relation to $f$-divergence
In the case of vanilla GAN (Goodfellow et al. 2014), the optimal discriminator is given by
@@ Appendix B Properties of MCGAN
Given (10), the vanilla GAN objective can be interpreted as minimizing Jensen-Shannon divergence between $\mu$ and $\nu_{\theta}$, subtracting a constant term $\log(4)$. The generator is therefore trained to minimize the Jensen-Shannon divergence. Similarly, (Mao et al. 2017) also showed that optimizing LSGANs yields minimizing the Pearson $\chi^{2}$ divergence between real and fake measures.
@@
Given (10), we are also able to explore the connection between MCGAN and $f$-divergence. As proven in Lemma 1, when considering the optimal discriminator (10), the difference between real and fake expected discriminator output in (6) can be interpreted as a $f$-divergence between $\mu$ and $\bar{\nu}$, where $\bar{\nu}:=\frac{(\mu+\nu)}{2}$ represents the averaged measure with density $p_{\bar{\nu}}(x):=\frac{p_{\mu}(x)+p_{\nu_{\theta}}(x)}{2}$.
Lemma 1.
@@
Lemma 1 establishes a connection between MCGAN and $f$-divergence, illustrating the information-theoretic aspects of the MCGAN framework. Unlike the $\text{KL}(\nu|\mu)$ induced in non-saturating loss (Arjovsky and Bottou 2017), this $\text{Div}_{f}(\mu|\bar{\nu})$ can avoid mode dropping by assigning moderate cost on the occasions where $p_{\mu}(x)\gg p_{\nu_{\theta}}(x)$.
Improved stability
Lack of stability is a well-known issue in GAN training, and it arises due to several factors. (Arjovsky and Bottou 2017) provides insights into this instability issue of non-saturating loss. The instability is analyzed by modeling the inaccurate discriminator as an optimal discriminator perturbed by a centered Gaussian process. Given this noisy version of the optimal discriminator, it can be shown that the gradient of non-saturating loss follows a centered Cauchy distribution with infinite mean and variance, which leads to massive and unpredictable updates of the generator parameter. Hence it can be regarded as the source of training instability.
@@
Theorem 2.
Let $D^{\phi_{\epsilon}}(x)$ be a noisy version of optimal discriminator such that $D^{\phi_{\epsilon}}(x)=D^{\phi^{*}}(x)+\epsilon_{1}(x)$ and $\nabla_{x}D^{\phi_{\epsilon}}(x)=\nabla_{x}D^{\phi^{*}}(x)+\epsilon_{2}(x)$ for $\forall x\in\mathcal{X},$ where $\epsilon_{1}(x)$ and $\epsilon_{2}(x)$ are two uncorrelated and centered random noises that are indexed by $x$ and have finite variance.333 This assumption of centered random noise is made due to the fact that as the approximation gets better, this error looks more and more like centered random noise due to the finite precision (Arjovsky and Bottou 2017). Then for $\mathcal{L}_{G}(\theta;\phi)$ in (3), we have
@@
Theorem 2 implies that given an inaccurate discriminator, the expected value of the gradient of $\mathcal{L}_{G}(\theta;\phi)$ in (3) is the accurate gradient given by the optimal discriminator and its variance is finite. More importantly, its variance is determined by the discrepancy between $\mu$ and $\nu_{\theta}$, specifically when the fake measure produces the real measure, the gradient is zero almost surely, indicating improved training stability.
Relation to feature matching
In order to enhance the generative performance, a feature matching approach is proposed in (Salimans et al. 2016) which adds to the generative loss function an additional cost that matches the statistic of the real and generated samples given by the activation on an intermediate layer of the discriminator. The generator hence is trained to generate fake samples that reflect the statistics (features maps) of real data rather than just maximizing its discriminator outputs.
@@
$R_{\text{fm}}(\theta;{\psi})=\|\mathbb{E}_{\mu}[\psi(x)]-\mathbb{E}_{\nu_{\theta}}[\psi(x)]\|^{2}_{2}.$ (13)